# Model-driven Stochastic Trace Clustering

### v1

@@ Appendix A Results for the Individual Event Logs
In this appendix we present the results for each event log separately. Columns correspond to the different metrics and rows with the different clustering techniques. We report weighted averages, i.e., to calculate the metric value for a clustering the measurements for each cluster are weighted by the amount of cases in that cluster.
@@
method replayfit replayprec alignfit alignprec simplarc ER graphdens graphentr
full log 0.989 0.781 0.978 0.781 0.682 3.618 0.279 12.729
random 0.992 0.695 0.984 0.695 0.690 2.754 0.268 7.912
frequencykm++ 0.995 0.831 0.990 0.831 0.686 2.670 0.259 4.779
trace2veckm++ 0.983 0.845 0.977 0.845 0.696 2.744 0.243 6.367
actitracdist 0.993 0.659 0.986 0.659 0.682 2.434 0.253 7.743
actitracfreq 0.919 0.885 0.918 0.886 0.686 3.026 0.266 8.807
EC++ 0.974 0.607 0.973 0.607 0.717 2.108 0.194 5.846
EC++norm 0.974 0.596 0.973 0.596 0.716 2.105 0.198 5.839
ECrand 0.969 0.664 0.968 0.664 0.699 2.118 0.216 6.094
ECsplit++ 0.967 0.750 0.954 0.750 0.702 2.379 0.276 7.241
ECsplit++norm 0.971 0.757 0.961 0.757 0.698 2.387 0.292 7.152
ECsplitrand 0.987 0.757 0.977 0.757 0.692 2.482 0.292 7.037
@@
method replayfit replayprec alignfit alignprec simplarc ER graphdens graphentr
full log 0.929 0.813 0.878 0.813 0.630 2.935 0.500 11.428
random 0.978 0.660 0.971 0.660 0.676 1.453 0.419 10.687
frequencykm++ 0.997 0.880 0.993 0.880 0.693 1.708 0.283 7.407
trace2veckm++ 0.991 0.925 0.987 0.925 0.746 1.229 0.341 9.038
actitracdist 0.994 0.825 0.992 0.825 0.768 0.961 0.369 5.157
actitracfreq 0.968 0.752 0.955 0.752 0.681 1.250 0.403 8.415
EC++ 0.995 0.818 0.988 0.818 0.776 0.730 0.333 8.567
EC++norm 0.946 0.682 0.950 0.682 0.726 0.730 0.300 7.720
ECrand 0.995 0.792 0.988 0.792 0.774 0.731 0.326 8.425
ECsplit++ 0.981 0.735 0.980 0.735 0.742 0.960 0.388 7.756
ECsplit++norm 0.981 0.687 0.980 0.687 0.676 0.960 0.366 7.780
ECsplitrand 0.981 0.665 0.980 0.665 0.707 0.958 0.355 7.733
@@
method replayfit replayprec alignfit alignprec simplarc ER graphdens graphentr
full log 0.990 0.591 0.975 0.591 0.629 3.450 0.416 18.223
random 0.993 0.463 0.979 0.463 0.639 2.307 0.318 15.437
frequencykm++ 0.995 0.637 0.987 0.637 0.648 2.156 0.258 11.383
trace2veckm++ 0.984 0.472 0.964 0.472 0.635 2.410 0.275 10.290
actitracdist 0.988 0.787 0.957 0.787 0.685 1.942 0.337 11.546
actitracfreq 0.991 0.627 0.988 0.627 0.631 1.968 0.304 12.834
EC++ 0.998 0.615 0.995 0.618 0.658 1.237 0.249 13.752
EC++norm 0.995 0.691 0.985 0.697 0.670 1.232 0.240 13.214
ECrand 0.995 0.674 0.984 0.677 0.671 1.239 0.239 13.512
ECsplit++ 0.988 0.669 0.967 0.669 0.666 1.356 0.309 14.327
ECsplit++norm 0.983 0.642 0.963 0.642 0.663 1.346 0.237 11.695
ECsplitrand 0.984 0.701 0.963 0.702 0.657 1.355 0.230 11.167
@@
method replayfit replayprec alignfit alignprec simplarc ER graphdens graphentr
full log 0.967 0.619 0.962 0.619 0.593 24.444 0.441 23.580
random 0.954 0.663 0.949 0.679 0.601 23.983 0.364 21.753
frequencykm++ 0.978 0.672 0.962 0.674 0.614 22.594 0.342 18.925
trace2veckm++ 0.972 0.602 0.956 0.602 0.629 23.175 0.499 18.212
actitracdist 0.968 0.567 0.967 0.575 0.606 22.695 0.367 20.646
actitracfreq 0.976 0.610 0.969 0.610 0.623 22.621 0.379 20.710
EC++ 0.961 0.532 0.948 0.536 0.601 21.163 0.326 18.841
EC++norm 0.974 0.539 0.958 0.543 0.603 21.324 0.327 18.977
ECrand 0.979 0.554 0.960 0.558 0.610 21.220 0.317 18.697
ECsplit++ 0.964 0.506 0.953 0.512 0.608 22.243 0.338 19.563
ECsplit++norm 0.971 0.486 0.959 0.490 0.596 22.039 0.325 19.497
ECsplitrand 0.986 0.455 0.971 0.457 0.612 22.040 0.332 19.332
@@
method replayfit replayprec alignfit alignprec simplarc ER graphdens graphentr
full log 0.966 0.841 0.957 0.841 0.733 10.728 0.533 3.668
random 0.959 0.866 0.953 0.866 0.734 10.498 0.586 3.558
frequencykm++ 0.987 0.947 0.930 0.960 0.770 9.704 0.522 3.079
trace2veckm++ 0.945 0.877 0.954 0.877 0.697 10.452 0.653 3.657
actitracdist 0.988 0.844 0.983 0.844 0.722 9.865 0.528 3.287
actitracfreq 0.988 0.775 0.983 0.775 0.722 9.867 0.516 3.305
EC++ 0.995 0.839 0.950 0.854 0.766 9.191 0.540 2.879
EC++norm 0.995 0.866 0.986 0.866 0.784 9.191 0.455 2.544
ECrand 0.991 0.833 0.979 0.833 0.771 9.445 0.513 2.953
ECsplit++ 0.997 0.909 0.964 0.915 0.769 9.767 0.537 3.319
ECsplit++norm 0.997 0.926 0.991 0.926 0.772 9.506 0.561 3.143
ECsplitrand 0.997 0.849 0.994 0.849 0.748 9.380 0.504 3.037
@@
method replayfit replayprec alignfit alignprec simplarc ER graphdens graphentr
full log 0.997 0.971 0.990 0.971 0.750 5.047 0.500 3.916
random 0.989 0.802 0.957 0.802 0.766 4.675 0.380 2.455
frequencykm++ 0.986 0.958 0.969 0.959 0.783 4.075 0.452 2.346
trace2veckm++ 0.993 0.941 0.977 0.941 0.776 4.717 0.448 2.711
actitracdist 0.987 0.901 0.970 0.901 0.775 4.342 0.453 2.819
actitracfreq 0.986 0.796 0.966 0.796 0.770 4.567 0.458 3.154
EC++ 0.997 0.907 0.989 0.907 0.772 3.826 0.371 1.622
EC++norm 0.996 0.813 0.989 0.813 0.774 3.792 0.337 1.625
ECrand 0.996 0.850 0.990 0.850 0.773 3.798 0.344 1.953
ECsplit++ 0.996 0.938 0.989 0.938 0.758 4.022 0.394 2.751
ECsplit++norm 0.998 0.813 0.994 0.813 0.768 3.921 0.407 2.782
ECsplitrand 0.996 0.922 0.990 0.922 0.764 3.975 0.399 2.588
@@
method replayfit replayprec alignfit alignprec simplarc ER graphdens graphentr
full log 0.999 0.150 / / 0.600 19.599 0.211 34.259
random 0.971 0.342 / / 0.625 18.803 0.204 32.639
frequencykm++ 0.973 0.532 / / 0.665 18.626 0.189 30.663
trace2veckm++ 0.971 0.535 / / 0.686 18.579 0.172 22.166
actitracdist 0.944 0.397 / / 0.652 17.326 0.191 27.323
actitracfreq 0.997 0.294 / / 0.643 16.963 0.176 26.933
EC++ 0.993 0.386 / / 0.640 16.798 0.171 24.821
EC++norm 0.993 0.369 / / 0.658 16.861 0.163 24.149
ECrand 0.964 0.680 / / 0.690 16.841 0.140 19.965
ECsplit++ 0.966 0.690 / / 0.714 17.030 0.151 20.643
ECsplit++norm 0.965 0.622 / / 0.699 16.980 0.151 20.009
ECsplitrand 0.969 0.624 / / 0.707 16.912 0.151 18.398
@@
method replayfit replayprec alignfit alignprec simplarc ER graphdens graphentr
full log 0.884 0.366 / / 0.572 26.273 0.308 29.932
random 0.901 0.415 / / 0.625 26.123 0.235 26.660
frequencykm++ 0.938 0.382 / / 0.600 24.584 0.235 24.045
trace2veckm++ 0.938 0.397 / / 0.612 24.351 0.246 22.264
actitracdist 0.905 0.401 / / 0.612 25.223 0.245 25.279
actitracfreq 0.930 0.364 / / 0.606 25.146 0.244 24.257
EC++ 0.937 0.397 / / 0.620 24.250 0.223 20.707
EC++norm 0.916 0.469 / / 0.625 24.387 0.240 22.087
ECrand 0.930 0.395 / / 0.623 24.330 0.239 22.198
ECsplit++ 0.943 0.354 / / 0.614 24.393 0.218 21.741
ECsplit++norm 0.940 0.432 / / 0.642 24.191 0.223 20.883
ECsplitrand 0.942 0.384 / / 0.616 24.249 0.217 20.916

### v2

@@ Appendix A Results for the Individual Event Logs
In this appendix, we present the results for each event log separately. Columns correspond to the different metrics and rows with the different clustering techniques. We report weighted averages, i.e., to calculate the metric value for a clustering, the measurements for each cluster are weighted by the number of cases in that cluster. Except for the column displaying ERsum which displays the total sum of the ER over all clusters (and all traces within those clusters). The weighted average ER is the same as the values in the ERsum column divided by the total number of cases in the event log. In addition, we provide figures for each cluster’s DFG in our online repository, together with the script used to generate them, which allows readers to visually compare the resulting models across clusters and methods.
@@
method replayfit replayprec alignfit alignprec simplarc ERav ERsum graphdens graphentr
full log 0.989 0.781 0.978 0.781 0.682 3.618 16570 0.279 12.729
random 0.992 0.695 0.984 0.695 0.690 2.754 12614 0.268 7.912
frequencykm++ 0.995 0.831 0.990 0.831 0.686 2.670 12230 0.259 4.779
trace2veckm++ 0.983 0.845 0.977 0.845 0.696 2.744 12566 0.243 6.367
actitracdist 0.993 0.659 0.986 0.659 0.682 2.434 11150 0.253 7.743
actitracfreq 0.919 0.885 0.918 0.886 0.686 3.026 13858 0.266 8.807
EC++ 0.974 0.607 0.973 0.607 0.717 2.108 9656 0.194 5.846
EC++norm 0.974 0.596 0.973 0.596 0.716 2.105 9641 0.198 5.839
ECrand 0.969 0.664 0.968 0.664 0.699 2.118 9702 0.216 6.094
ECsplit++ 0.967 0.750 0.954 0.750 0.702 2.379 10894 0.276 7.241
ECsplit++norm 0.971 0.757 0.961 0.757 0.698 2.387 10930 0.292 7.152
ECsplitrand 0.987 0.757 0.977 0.757 0.692 2.482 11369 0.292 7.037
@@
method replayfit replayprec alignfit alignprec simplarc ERav ERsum graphdens graphentr
full log 0.929 0.813 0.878 0.813 0.630 2.935 441317 0.500 11.428
random 0.978 0.660 0.971 0.660 0.676 1.453 218421 0.419 10.687
frequencykm++ 0.997 0.880 0.993 0.880 0.693 1.708 256904 0.283 7.407
trace2veckm++ 0.991 0.925 0.987 0.925 0.746 1.229 184810 0.341 9.038
actitracdist 0.994 0.825 0.992 0.825 0.768 0.961 144564 0.369 5.157
actitracfreq 0.968 0.752 0.955 0.752 0.681 1.250 187920 0.403 8.415
EC++ 0.995 0.818 0.988 0.818 0.776 0.730 109777 0.333 8.567
EC++norm 0.946 0.682 0.950 0.682 0.726 0.730 109721 0.300 7.720
ECrand 0.995 0.792 0.988 0.792 0.774 0.731 109866 0.326 8.425
ECsplit++ 0.981 0.735 0.980 0.735 0.742 0.960 144290 0.388 7.756
ECsplit++norm 0.981 0.687 0.980 0.687 0.676 0.960 144333 0.366 7.780
ECsplitrand 0.981 0.665 0.980 0.665 0.707 0.958 144564 0.355 7.733
@@
method replayfit replayprec alignfit alignprec simplarc ERav ERsum graphdens graphentr
full log 0.990 0.591 0.975 0.591 0.629 3.450 345039 0.416 18.223
random 0.993 0.463 0.979 0.463 0.639 2.307 230692 0.318 15.437
frequencykm++ 0.995 0.637 0.987 0.637 0.648 2.156 215589 0.258 11.383
trace2veckm++ 0.984 0.472 0.964 0.472 0.635 2.410 241045 0.275 10.290
actitracdist 0.988 0.787 0.957 0.787 0.685 1.942 194174 0.337 11.546
actitracfreq 0.991 0.627 0.988 0.627 0.631 1.968 196769 0.304 12.834
EC++ 0.998 0.615 0.995 0.618 0.658 1.237 123719 0.249 13.752
EC++norm 0.995 0.691 0.985 0.697 0.670 1.232 123175 0.240 13.214
ECrand 0.995 0.674 0.984 0.677 0.671 1.239 123922 0.239 13.512
ECsplit++ 0.988 0.669 0.967 0.669 0.666 1.356 135583 0.309 14.327
ECsplit++norm 0.983 0.642 0.963 0.642 0.663 1.346 134598 0.237 11.695
ECsplitrand 0.984 0.701 0.963 0.702 0.657 1.355 135463 0.230 11.167
@@
method replayfit replayprec alignfit alignprec simplarc ERav ERsum graphdens graphentr
full log 0.967 0.619 0.962 0.619 0.593 24.444 25666 0.441 23.580
random 0.954 0.663 0.949 0.679 0.601 23.983 25182 0.364 21.753
frequencykm++ 0.978 0.672 0.962 0.674 0.614 22.594 23724 0.342 18.925
trace2veckm++ 0.972 0.602 0.956 0.602 0.629 23.175 24334 0.499 18.212
actitracdist 0.968 0.567 0.967 0.575 0.606 22.695 23829 0.367 20.646
actitracfreq 0.976 0.610 0.969 0.610 0.623 22.621 23752 0.379 20.710
EC++ 0.961 0.532 0.948 0.536 0.601 21.163 22221 0.326 18.841
EC++norm 0.974 0.539 0.958 0.543 0.603 21.324 22391 0.327 18.977
ECrand 0.979 0.554 0.960 0.558 0.610 21.220 22281 0.317 18.697
ECsplit++ 0.964 0.506 0.953 0.512 0.608 22.243 23355 0.338 19.563
ECsplit++norm 0.971 0.486 0.959 0.490 0.596 22.039 23141 0.325 19.497
ECsplitrand 0.986 0.455 0.971 0.457 0.612 22.040 23142 0.332 19.332
@@
method replayfit replayprec alignfit alignprec simplarc ERav ERsum graphdens graphentr
full log 0.966 0.841 0.957 0.841 0.733 10.728 81036 0.533 3.668
random 0.959 0.866 0.953 0.866 0.734 10.498 79305 0.586 3.558
frequencykm++ 0.987 0.947 0.930 0.960 0.770 9.704 73301 0.522 3.079
trace2veckm++ 0.945 0.877 0.954 0.877 0.697 10.452 78958 0.653 3.657
actitracdist 0.988 0.844 0.983 0.844 0.722 9.865 74520 0.528 3.287
actitracfreq 0.988 0.775 0.983 0.775 0.722 9.867 74536 0.516 3.305
EC++ 0.995 0.839 0.950 0.854 0.766 9.191 69431 0.540 2.879
EC++norm 0.995 0.866 0.986 0.866 0.784 9.191 69429 0.455 2.544
ECrand 0.991 0.833 0.979 0.833 0.771 9.445 71346 0.513 2.953
ECsplit++ 0.997 0.909 0.964 0.915 0.769 9.767 73778 0.537 3.319
ECsplit++norm 0.997 0.926 0.991 0.926 0.772 9.506 71809 0.561 3.143
ECsplitrand 0.997 0.849 0.994 0.849 0.748 9.380 70854 0.504 3.037
@@
method replayfit replayprec alignfit alignprec simplarc ERav ERsum graphdens graphentr
full log 0.997 0.971 0.990 0.971 0.750 5.047 7505 0.500 3.916
random 0.989 0.802 0.957 0.802 0.766 4.675 6952 0.380 2.455
frequencykm++ 0.986 0.958 0.969 0.959 0.783 4.075 6060 0.452 2.346
trace2veckm++ 0.993 0.941 0.977 0.941 0.776 4.717 7014 0.448 2.711
actitracdist 0.987 0.901 0.970 0.901 0.775 4.342 6457 0.453 2.819
actitracfreq 0.986 0.796 0.966 0.796 0.770 4.567 6791 0.458 3.154
EC++ 0.997 0.907 0.989 0.907 0.772 3.826 5690 0.371 1.622
EC++norm 0.996 0.813 0.989 0.813 0.774 3.792 5639 0.337 1.625
ECrand 0.996 0.850 0.990 0.850 0.773 3.798 5648 0.344 1.953
ECsplit++ 0.996 0.938 0.989 0.938 0.758 4.022 5981 0.394 2.751
ECsplit++norm 0.998 0.813 0.994 0.813 0.768 3.921 5830 0.407 2.782
ECsplitrand 0.996 0.922 0.990 0.922 0.764 3.975 5911 0.399 2.588
@@
method replayfit replayprec alignfit alignprec simplarc ERav ERsum graphdens graphentr
full log 0.999 0.150 / / 0.600 19.599 256496 0.211 34.259
random 0.971 0.342 / / 0.625 18.803 246081 0.204 32.639
frequencykm++ 0.973 0.532 / / 0.665 18.626 243757 0.189 30.663
trace2veckm++ 0.971 0.535 / / 0.686 18.579 243138 0.172 22.166
actitracdist 0.944 0.397 / / 0.652 17.326 226741 0.191 27.323
actitracfreq 0.997 0.294 / / 0.643 16.963 221992 0.176 26.933
EC++ 0.993 0.386 / / 0.640 16.798 219834 0.171 24.821
EC++norm 0.993 0.369 / / 0.658 16.861 220660 0.163 24.149
ECrand 0.964 0.680 / / 0.690 16.841 220403 0.140 19.965
ECsplit++ 0.966 0.690 / / 0.714 17.030 222875 0.151 20.643
ECsplit++norm 0.965 0.622 / / 0.699 16.980 222212 0.151 20.009
ECsplitrand 0.969 0.624 / / 0.707 16.912 221330 0.151 18.398
@@
method replayfit replayprec alignfit alignprec simplarc ERav ERsum graphdens graphentr
full log 0.884 0.366 / / 0.572 26.273 148365 0.308 29.932
random 0.901 0.415 / / 0.625 26.123 147466 0.235 26.660
frequencykm++ 0.938 0.382 / / 0.600 24.584 138775 0.235 24.045
trace2veckm++ 0.938 0.397 / / 0.612 24.351 137463 0.246 22.264
actitracdist 0.905 0.401 / / 0.612 25.223 142486 0.245 25.279
actitracfreq 0.930 0.364 / / 0.606 25.146 142026 0.244 24.257
EC++ 0.937 0.397 / / 0.620 24.250 136892 0.223 20.707
EC++norm 0.916 0.469 / / 0.625 24.387 137666 0.240 22.087
ECrand 0.930 0.395 / / 0.623 24.330 137342 0.239 22.198
ECsplit++ 0.943 0.354 / / 0.614 24.393 137698 0.218 21.741
ECsplit++norm 0.940 0.432 / / 0.642 24.191 136560 0.223 20.883
ECsplitrand 0.942 0.384 / / 0.616 24.249 136884 0.217 20.916
Appendix B Cluster count selection (Elbow Analysis)
This appendix presents the detailed results of the parameter determination phase. To select an appropriate number of clusters ($k$) for each event log, we analyzed the evolution of model quality metrics across a range of cluster counts, from $k=1$ (the original unclustered log) to $k=10$. For each $k$, we executed four clustering variants: random clustering, frequency-based clustering (with kmeans++), EC, and ECsplit (both utilizing ++ initialization). To ensure a representative evaluation of the resulting partition, the metric scores reported in the figures below are calculated as the weighted average over the generated clusters, where the weight corresponds to the number of traces assigned to each cluster. The selection of $k$ was based on the “elbow method”, identifying the point where the marginal improvement in the metrics (e.g., reduction in Entropic Relevance) diminishes significantly. Figure 5 displays the ER (average per trace) scores for different numbers of clusters, Figure 6 the graph density, and Figure 7 the graph entropy. Scores for $k=1$ correspond to no clustering, i.e. calculated on the whole event log.
Figure 5: The elbow experiment with (simplified) ER, averaged per trace.
Figure 6: The elbow experiment with graph density.
Figure 7: The elbow experiment with graph entropy.